OpenStudy (anonymous):

The first five terms of a sequence are given.6,2,2/3,2/9,2/27. What equation describes the nth term of the sequence?

OpenStudy (ivettef365):

what is the sequence ??

OpenStudy (anonymous):

6, 2, 2/3, 2/9, and 2/27

OpenStudy (ivettef365):

do you have choices

OpenStudy (anonymous):

yeah, t(n)=6n t(n)=2*(3)^n t(n)=18(1/3)^n t(n)=16(1/3)n

OpenStudy (ivettef365):

ok the one that makes sense is the 3rd choice 18(1/3)^n

OpenStudy (ivettef365):

if you look at the numbers the sequence goes by multiplying by 1/3 6 x 1/3 = 2 2 x 1/3 = 2/3 2/3 x 1/3 = 2/9 2/9 x 1/3 = 2/27 and so on

OpenStudy (anonymous):

thanks!

OpenStudy (ivettef365):

yw :)